# Unscented Kalman Filter with a Nonlinear Propagation Model for Navigation Applications

### v1

@@ I Introduction
Regardless if the filter is adaptive or not, the UKF uses a dynamic model to propagate the state vector and the covariance matrix. Having an inaccurate dynamic model can decrease performance and may lead to filter divergence. When applying the UKF to improve the performance of a navigation system, some authors used the navigation solution as the state vector as in [9]. It is also common to use the navigation error solution as the state vector, enhancing the accuracy of the unscented transform [10]. There, a model predictive UKF was introduced for inertial navigation system/global navigation satellite system (INS/GNSS) fusion.
In this paper, we introduce the unscented Kalman filter navigation error state propagating method (UKF-NESPM) to accurately propagate the sigma points using the navigation algorithm itself, enabling us to capture the exact dynamic functionalities. We demonstrate the benefits of our proposed approach using real sensor data recorded by an autonomous underwater vehicle during several scenarios.
@@ II-A The UKF algorithm
    where $\boldsymbol{\mathbf{P}}_{{k}{|}{k}}$ is positive definite, therefore it can be factored by Cholesky decomposition as $\sqrt{\boldsymbol{\mathbf{P}}_{{k}{|}{k}}}\sqrt{\boldsymbol{\mathbf{P}}_{{k}{|}{k}}}^{T}$. $\sqrt{\boldsymbol{\mathbf{P}}_{{k}{|}{k}}}_{i}$ is the ${i}$ row of the matrix $\sqrt{\boldsymbol{\mathbf{P}}_{{k}{|}{k}}}$. The points are symmetrical around the expected value, so the expectation is preserved. ${\lambda=\alpha^{2}(n+\kappa)-n}$ is a scaling parameter, $\alpha$ determines the spread of the sigma points around $\hat{{x}}_{{k}{|}{k}}$, usually set to a small value ($31e-$ 3), $\kappa$ is a secondary scaling parameter usually set to 0, and $\beta$ is used to incorporate prior knowledge of the distribution of $x$ (for Gaussian distribution it is set to 2).
@@
    $\centering\boldsymbol{x}_{i,{{k+1}{|}{k}}}=f(\boldsymbol{x}_{i,{{k}{|}{k}}}),i=0,...,2n\@add@centering$ (11)
@@
    where $\delta\boldsymbol{x}_{i,{{k+1}{|}{k}}}=\boldsymbol{x}_{i,{{k+1}{|}{k}}}-\hat{\boldsymbol{x}}_{{k+1}{|}{k}}$, $\boldsymbol{\mathbf{Q}}_{k+1}$ is the covariance process noise matrix at step ${k+1}$
@@ II-B The Inertial Navigation Algorithm
    Where T is the transform operator, $dt$ is the cycle duration, and
    ${w^{*}}_{nb,k+1}^{b}=w_{ib,k+1}^{b}-C_{n,k}^{b}(w_{ie,k+1}^{n}+w_{en,k+1}^{n})$ (25)
    where $w_{ib,k+1}^{b}$ is the angular velocity of the body frame relate to the inertial system, resolved in the body frame, during time step $k+1$ in the body frame, given by the gyroscopes, $w_{ie,k+1}^{n}$ is the earth rotation relate to the inertial frame, given in the navigation frame computed at time step $k+1$, and $w_{en,k+1}^{n}$ is the angular velocity of the navigation frame relate to earth resolved in the navigation frame computed at time step $k+1$.
@@ III Proposed Method
where $\xi\in\mathbb{R}^{3\times 1}$ is the position vector consisting of the latitude, longitude, and altitude, $v^{n}\in\mathbb{R}^{3\times 1}$ is the velocity error expressed in the navigation frame, and, $\Psi^{n}\in\mathbb{R}^{3\times 1}$ is the Euler angles representing the transformation from the navigation frame to the body frame. Let ${\sigma}_{k}\in\mathbb{R}^{12\times 1}$ be a given sigma point at time step $k$ such that:
@@
$\centering{{x_{\sigma}}^{\psi}}_{k}={{x_{m}}^{\psi}}_{k}\oplus{{\sigma}^{\delta\psi}}_{k}\@add@centering$ (37)
@@
where ${{f^{*}}^{b}}_{k+1}$ and ${{w^{*}}^{b}}_{k+1}$ are the inertial measurements compensated by the appropriate sigma point biases. To extract the propagated sigma point, we need to subtract from each solution the propagated mean solution, resulting with the propagated sigma points:
@@ IV-A Dataset
We used the Snapir AUV dataset [12] for training and testing our proposed approach. Snapir is an ECA Robotics modified Group A18D mid-size AUV. The Snapir AUV is outfitted with the iXblue, Phins Subsea INS, which uses fiber optic gyroscope for precise inertial navigation [13]. Snapir also uses a Teledyne RDI Work Horse Navigator Doppler velocity log (DVL) [14], known for its capability to provide accurate velocity measurements. The INS operates at a frequency of 100 [Hz], whereas the DVL operates at 1[Hz]. The dataset contains 24 minutes of recording divided into 6 trajectories. Each trajectory contains the DVL/INS fusion solution, the ground truth (GT), inertial readings, and DVL measurements. To create the trajectories being tested, we added the following to the given raw data: initial velocity errors with a standard deviation (STD) of $0.25[m/s]$ and $0.05[m/s]$, horizontal and vertical, respectively, and misalignment errors with a STD of $0.01[deg]$ per axis. We also added noises with a STD of $0.03[m/s^{2}]$ to the accelerometers and $7.3\times 10^{-6}[rad/s]$ to the gyroscopes. Finally, we added biases with a STD of $0.3[m/s^{2}]$ to the accelerometers and $7.3\times 10^{-5}[rad/s]$ to the gyroscopes.
For the training process, we used tracks 1-4, each consisting of 4 minutes, resulting in a total of 16 minutes of the training dataset. The trajectories, presented in Figure 1 include a wide range of dynamics and maneuvers. Track 1 (Figure 1(a)) and track 2 (Figure 1(b)) have relatively more maneuvers and dynamic changes than track 3 (Figure 1(c)) and track 4 (Figure 1(d)), has longer straight sections.
@@
(c) Track 3
(d) Track 4
Figure 1: Horizontal position of tracks 1-4 used in the training dataset.
@@ IV-C Results
We applied the two filters, ANUKF and ANUKF-NESPM, to the test dataset. Table I summarizes the VRMSE and MRMSE results of the two filters, showing that the velocity solution of the ANUKF-NESPM performs better than the ANUKF. In other words, ANUKF-NESPM showed an improvement of $8.8\%$ over ANUKF. The ANUKF-NESPM estimation of misalignment showed an improvement of $17.7\%$ over ANUKF.
@@ V Conclusion
To demonstrate the performance of our approach, we focused on the INS/DVL fusion problem for a maneuvering AUV. We evaluated our approach compared to the adaptive neural UKF using a real-world recorded AUV dataset. In the VRMSE metric, ANUKF-NESPM improved by $8.8\%$ over ANUKF. In the MRMSE metric, ANUKF-NESPM improved by $17.7\%$ over ANUKF. In conclusion our ANUKF-NESPM offers an accurate navigation solution in normal operating conditions of INS/DVL fusion. Thus, it allows planning and operating in challenging AUV tasks, including in varying dynamics.

### v2

@@ I Introduction
Regardless if the filter is adaptive or not, the UKF uses a dynamic model to propagate the state vector and the covariance matrix. Having an inaccurate dynamic model can decrease performance and may lead to filter divergence. When applying the UKF to improve the performance of a navigation system, some authors used the navigation solution as the state vector as in [9] and [10]. It is also common to use the navigation error solution as the state vector, enhancing the accuracy of the unscented transform [11]. There, a model predictive UKF was introduced for inertial navigation system/global navigation satellite system (INS/GNSS) fusion.
In this paper, we introduce the unscented Kalman filter navigation error state propagating method (UKF-NESPM) to accurately propagate the sigma points using the navigation algorithm itself, enabling us to capture the exact dynamic functionalities. We demonstrate the benefits of our proposed approach using real sensor data recorded by an autonomous underwater vehicle (AUV) during several scenarios.
@@ II-A The UKF algorithm
[…]
@@
    $\centering\boldsymbol{x}_{i,{{k+1}{|}{k}}}=f(\boldsymbol{x}_{i,{{k+1}{|}{k}}}),i=0,...,2n\@add@centering$ (11)
@@
    where $(\circ)^{T}$ is the transform operator, $\delta\boldsymbol{x}_{i,{{k+1}{|}{k}}}=\boldsymbol{x}_{i,{{k+1}{|}{k}}}-\hat{\boldsymbol{x}}_{{k+1}{|}{k}}$, $\boldsymbol{\mathbf{Q}}_{k+1}$ is the covariance process noise matrix at step ${k+1}$
@@ II-B The Inertial Navigation Algorithm
    where $dt$ is the cycle duration ($\frac{dt}{2}$ was used to calculate the average attitude during step $k+1$ to project the acceleration measurements from body frame to the navigation frame), and
    ${w^{*}}_{nb,k+1}^{b}=w_{ib,k+1}^{b}-C_{n,k}^{b}(w_{ie,k}^{n}+w_{en,k}^{n})$ (25)
    where $w_{ib,k+1}^{b}$ is the angular velocity of the body frame relate to the inertial system, resolved in the body frame, during time step $k+1$ in the body frame, given by the gyroscopes, $w_{ie,k}^{n}$ is the earth rotation relate to the inertial frame, given in the navigation frame computed at time step $k$, and $w_{en,k}^{n}$ is the angular velocity of the navigation frame relate to earth resolved in the navigation frame computed at time step $k$. The symbol * denotes intermediate calculation phase.
@@ III Proposed Method
where $\xi\in\mathbb{R}^{3\times 1}$ is the position vector consisting of the latitude, longitude, and altitude, $v^{n}\in\mathbb{R}^{3\times 1}$ is the velocity expressed in the navigation frame, and, $\Psi^{n}\in\mathbb{R}^{3\times 1}$ is the Euler angles representing the transformation from the navigation frame to the body frame. Let ${\sigma}_{k}\in\mathbb{R}^{12\times 1}$ be a given sigma point (error state vector) at time step $k$ such that:
@@
$\centering{{x_{\sigma}}^{\psi}}_{k}={{x}^{\psi}}_{k}\oplus{{\sigma}^{\delta\psi}}_{k}\@add@centering$ (37)
@@
where ${{f^{*}}^{b}}_{k+1}$ and ${{w^{*}}^{b}}_{k+1}$ are the inertial measurements compensated by the appropriate sigma point biases. To extract the propagated sigma point, we need to subtract from each solution the propagated mean solution, denoted as $x_{m}$, resulting with the propagated sigma points:
@@ IV-A Dataset
We used the Snapir AUV (Figure 1) dataset [13] for training and testing our proposed approach. Snapir is an ECA Robotics modified Group A18D mid-size AUV. The Snapir AUV is outfitted with the iXblue, Phins Subsea INS, which uses fiber optic gyroscope for precise inertial navigation [14]. Snapir also uses a Teledyne RDI Work Horse Navigator Doppler velocity log (DVL) [15], known for its capability to provide accurate velocity measurements. The INS operates at a frequency of 100 [Hz], whereas the DVL operates at 1[Hz].
Figure 1: The Snapir AUV during sea experiments.
[…]
For the training process, we used tracks 1-4, each consisting of 4 minutes, resulting in a total of 16 minutes of the training dataset. The trajectories presented in Figure 2 and Figure 3 include a wide range of dynamics and maneuvers. Track 1 (Figure 2(a)) and track 2 (Figure 2(b)) have relatively more maneuvers and dynamic changes than track 3 (Figure 3(a)) and track 4 (Figure 3(b)), has longer straight sections.
@@
Figure 2: Horizontal position of tracks 1-2 used in the training dataset.
(a) Track 3
(b) Track 4
Figure 3: Horizontal position of tracks 3-4 used in the training dataset.
@@ IV-C Results
[…]
@@ V Conclusion
[…]
Acknowledgements
The authors gratefully acknowledge the support of the Israel Innovation Authority under grant 84330, which partially supported this research.